# Fluid mechanics, Applied mathematics, Mathematical modelling

### v2

@@ 3.1 Local dynamics
We begin by considering the model’s local dynamics – the homogeneous dynamics in the absence of spatial derivatives. One may naturally think of these as the dynamics of horizontally uniform states with no $(x,z)$ dependence, such as fully laminar flow or fully turbulent flow, or one may think more loosely of these as the model dynamics within a small domain, a minimal flow unit [49, 40], in which large-scale spatial variation is not possible. From either perspective, the local dynamics are very informative of the core model behaviour.
Dropping all spatial derivatives in the full model equations, one immediately sees that the only fields that play a role in the local dynamics are $u_{1}$ and $q_{0}$. The other fields decouple and are either necessarily zero by incompressibility, as is the case for $v_{1}$, or they undergo exponential decay to zero, as in the case of $u_{0},w_{0}$, and $w_{1}$. We let $u_{1}(x,z,t)\rightarrow u(t)$ and $q_{0}(x,z,t)\rightarrow q(t)$. Then the local dynamics are governed by the ordinary differential equations
@@
These equations capture the essence of the interaction between the sinusoidal shear profile with amplitude $u$ and turbulent kinetic energy with amplitude $q$. They contain the two most important model closures: $A(q)$ modelling the Reynolds-stress amplitude and $\varepsilon_{0}(q;Re)$ modelling the dominant component of the turbulent pseudo-dissipation rate. (See equations (11), (14), and (15).) The $1$ appearing in the expression $(1-u)$ in momentum equation (20a) comes from the body force that acts against viscous decay and Ekman friction to drive $u$ to toward the laminar value $u=1$. The Reynolds-stress divergence, $-\beta A(q)$, is a force acting against the shear whenever $q>0$. Recall that $A(0)=0$. In the TKE equation (20b), TKE is generated by production, $\frac{1}{2}\beta uA(q)$, and dissipated by Ekman friction and the turbulent pseudo-dissipation rate.
The local dynamics described by equations (20) can be understood in terms of nullclines: curves in the $(u,q)$ phase plane on which $\dot{u}=0$ and $\dot{q}=0$. These are shown in figure 6a for two values of $Re$. Steady states of the local dynamics occur where the nullclines intersect and hence $\dot{u}=\dot{q}=0$. We denote steady states by $(u_{ss},q_{ss})$. One steady state is laminar flow $(u_{ss},q_{ss})=(1,0)$ that exists for all $Re$. For $Re$ above $Re_{\rm sn}=72.4$, the nullclines intersect at a pair of steady states with $u_{ss}<1$ and $q_{ss}>0$, corresponding to turbulent states. These states are shown in the bifurcation diagram in figure 6b, with turbulent states referred to as upper branch and lower branch depending on the value of $q_{ss}$. These local dynamics are very similar to those in models exhibiting a Turing instability [18, 38].
@@
Figure 6: (a) Nullclines and (b) bifurcation diagram for local dynamics. In (a), the $u$ nullclines, on which $\dot{u}=0$, are shown in blue and the $q$ nullclines, on which $\dot{q}=0$, are shown in red. Nullclines intersect at steady states of the local dynamics. At $Re=65$, the only steady state is laminar flow $(u_{ss},q_{ss})=(1,0)$, while at $Re=85$, there are two additional steady states. In (b), steady states are plotted as a function of $Re$. $Re_{\rm sn}=72.4$ marks a saddle-node bifurcation in which upper- and lower-branch solutions emerge. Lower-branch solutions are unstable; upper branch solutions are unstable to patterns below $Re_{c}=85.1$ and stable above.
@@ 3.2 Turbulent bands
Going beyond the local dynamics, the model reproduces many features of turbulent bands observed in experiments and DNS of planar shear flows: plane Couette flow, Waleffe flow, and plane Poiseuille flow. We illustrate here some of the most significant dynamics captured by the model. For this we simulate the full model equations on a doubly-periodic domain of size $L_{x}\times L_{z}$ using the pseudo-spectral code Dedalus [44]. Second-order Runge-Kutta time stepping is used with a time step $\Delta t=0.04$, approximately 10 times larger than can be used in a corresponding DNS of the Navier-Stokes equations. A Fourier-spectral method with $3/2$ dealiasing is used with a resolution of one grid point per space unit, approximately ten times fewer than for a DNS. Based on the numerical time step and $(x,z)$ spatial discretizations alone, the model is $O(10^{3})$ faster to simulate than the Navier-Stokes equations. Two types of initial conditions (ICs) are used to initialize simulations. A localized IC comprises a finite region of turbulence of length $25$ and width $14$ tilted at $24^{\circ}$ to the $x$ direction. The second IC is uniform turbulence $(u_{0},u_{1},v_{1},w_{0},w_{1},q_{0})(x,z,0)=(0,u_{ss},0,0,0,q_{ss})$. In all cases, the initial large-scale flow fields are additionally seeded with small-scale noise. (To account for turbulent fluctuations on the “microscopic scale”, noise may be further included in simulations. We do not consider this here. The only source of randomness is the ICs.)
@@
Figure 7a shows a simulation at $Re=75$ starting from the localized IC. During the initial evolution, a large-scale, quadrupolar velocity field is established. Such fields are well documented [50, 51, 36, 25, 34, 46, 52, 53, 54], and reproducing them is an important validation of the model. The turbulent patch elongates through the growth of its two tips. They eventually join and the system settles into a single steady, straight turbulent band tilted at $\theta=24^{\circ}$ to the streamwise direction [55, Movies S1-S2]. While the final angle of the band to the streamwise direction is constrained by the domain, during the initial band elongation (figure 7a, $t=750$), the angle is not far from $\theta=24^{\circ}$ (see also figure 9a, $t=2700$, discussed below). Furthermore, doubling the domain size in the spanwise direction still results in $24^{\circ}$ bands [55, Movie S8]. This is consistent with numerous observations of bands where angles typically line in the range $24^{\circ}\lesssim\theta\lesssim 45^{\circ}$ [4, 12, 22, 24, 25, 56, 19, 13, 57, 58]. The large-scale flow directed along the edges of the turbulent bands is typical of that observed in experiments and DNS [48, 7, 25, 9, 59]. It is worth emphasizing that the large-scale velocity vectors seen in the midplane are given by $u_{0}(x,z){\bf e}_{x}+w_{0}(x,z){\bf e}_{z}$.
@@
Figure 7b shows a simulation at $Re=78$. At this slightly higher value of $Re$, we observe more growth at the tips and also lateral splitting of the band, eventually resulting in two identical steady bands within the domain [55, Movies S3-S4]. Figure 7c shows the $Re=78$ case again, but now starting from the uniform IC. Initially, there is competition between symmetrically related orientations (the system is symmetric under spanwise reflections, $z\to-z$). Eventually, one orientation emerges and a single pair of steady bands forms [55, Movie S5]. Both lateral splitting [23, 24, 60, 59, 61] and band competition [15, 22, 24, 12, 13, 62, 19] are well-documented phenomena.
@@
Figure 7d shows a vertical slice of the reconstructed turbulence field $q$ (top panel) and large-scale velocities (bottom two panels), taken across the final steady band at $Re=75$ (cyan line in figure 7a, t = 25350). The plots are strikingly similar to vertical slices in DNS of turbulent bands [7, 34, 9, 4]. The top panel highlights the so-called overhang turbulent regions [48, 7, 25] that are generated through advection of the turbulent field by the mean shear and captured by the $q_{1}$ field in the model. The middle panel visualizes the streamlines of the large-scale flow in the slice plane and the bottom panel shows contours of the well-documented along-band flow (through the slice plane). The structure in the wall-normal direction seen in figure 7d highlights that, while model fields are functions of $(x,z)$, they describe non-trivial, three-dimensional flows very much like those seen in DNS.
@@
We now survey the spatiotemporal dynamics of bands at a fixed angle and independent of the along-band direction. Restricting to such solutions, the model reduces to a system of 1D PDEs in the across-band coordinate. While restricting the band angle eliminates fully two-dimensional phenomena, such as seen in the early times in figure 7, it provides insights into some of the key spatiotemporal features of turbulent bands.
@@
Figure 8 shows spatiotemporal diagrams at representative values of $Re$, with the band angle fixed at $\theta=24^{\circ}$. Simulations are initiated from a localized IC (a steady band solution at $Re=66$), except in figure 8f where the simulation is initiated with a uniform turbulent IC. At sufficiently low $Re$, turbulent bands are not sustainable and localized initial states rapidly decay, as illustrated at $Re=64$. Note the timescale of panel figure 8a in comparison with the other panels. Starting at $Re\simeq 66$, steady bands are sustained in the form of localized, isolated structures. This is well below the onset of bistability in the homogeneous dynamics in figure 6.
@@
At larger $Re$, figures 8c-8e, an initially localized turbulent patch spreads laterally. At $Re=80$ and $Re=82$, the spreading occurs via what is known as splitting, and the resulting asymptotic states are regular periodic patterns of turbulent bands. (Beyond time shown in figure 8c, the system reaches a steady pattern with eight turbulent bands.) At $Re=90$, turbulence expands rapidly in what is referred to as a slug [6, 63, 28]. The asymptotic state is uniform turbulence. As can be seen, the expansion rate is an increasing function of $Re$. The final panel, figure 8e, shows the formation of bands from uniform turbulence at $Re=80$, the same $Re$ as in figure 8c. This demonstrates explicitly the instability of the uniform turbulent state at this $Re$.
@@
Figure 8: Spatiotemporal dynamics of 1D bands at a fixed angle $\theta=24^{\circ}$, for various $Re$ indicated in the panels. Visualized is the turbulent field $q_{0}(z^{\prime},t)$, where $z^{\prime}$ is the across-band coordinate. (a)-(e) Dynamics starting from a localized turbulent patch illustrating the following behaviour: (a) relaminarization, (b) single, localized turbulent band, (c) and (d) expanding turbulence in the form of band splitting or patterned slug, (e) expanding turbulent slug. (f) Illustration of band formation from uniform turbulence. Panels (c) and (f) correspond to the same conditions, just different initial conditions. Beyond the time shown in (c) the system develops eight turbulent bands.
@@
Finally, returning to the full 2D model dynamics, we show in figure 9 some of the model’s behaviour in larger domains—twice the size in each horizontal dimension compared to figure 7—starting from the localized ICs. While the initial evolution (first column of figure 9) resembles that of runs in figure 7, the larger domain better captures the complex process of turbulence proliferation. The single band in figure 9a elongates through the growth of its two tips, as before, but no longer simply reconnects with itself upon reaching the boundaries. Instead, it produces a complex web of chaotic bands [55, Movie S9]. At slightly higher $Re$, proliferation changes markedly due to band splitting (figure 9b). Turbulence expands in all directions—albeit anisotropically—via the continuous generation of new bands through splitting [55, Movie S10]. This spreading behaviour is strikingly similar to that observed in plane Couette flow [24, 59]. At yet higher Reynolds number (figure 9c), the initial spreading occurs more like a front of nearly uniform turbulence, which subsequently develops laminar gaps, resulting in bands [55, Movie S11]. For all cases, intermediate times reveal long-lived complex dynamics as the system adjusts toward equilibrium. During this phase, bands of different orientations compete (middle column of figure 9). Not all combinations of ICs, Reynolds number, and domain size lead to simple bands. We observe a variety of other steady states, such as criss-crossing bands [55, Movie S6], non-uniform bands (figure 9b), and instances featuring simple bands of both orientations separated by domain boundaries (figure 9c) [15, 22]. Some runs never reach a steady state: bands nearly form, become unstable, break up, and re-form in a repeating cycle. The dynamics of the model are rich, and there is much more to explore, but we leave this for future work.
@@
Figure 9: Illustrative model dynamics in large domains. Temporal evolution of turbulent bands for three values of $Re$, all initialized with the same localized initial condition. (a) $Re=75$, (b) $Re=78$, and (c) $Re=80$ [55, Movies S9-S11, respectively]. The colour represents the amplitude of $q_{0}$, the vertically-averaged turbulent kinetic energy.

### v3

@@ 3.1 Local dynamics
We begin by considering the model’s local dynamics – the homogeneous dynamics in the absence of spatial derivatives. One may naturally think of these as the dynamics of horizontally uniform states with no $(x,z)$ dependence, such as fully laminar flow or fully turbulent flow, or one may think more loosely of these as the model dynamics within a small domain, a minimal flow unit [hamilton1995regeneration, waleffe1997self], in which large-scale spatial variation is not possible. From either perspective, the local dynamics are very informative of the core model behaviour.
Dropping all spatial derivatives in the model equations, one immediately sees that the only fields that play a role in the local dynamics are $u_{1}$ and $q_{0}$. The other fields decouple and are either necessarily zero by incompressibility, as is the case for $v_{1}$, or they undergo exponential decay to zero, as in the case of $u_{0},w_{0}$, and $w_{1}$. We let $u_{1}(x,z,t)\rightarrow u(t)$ and $q_{0}(x,z,t)\rightarrow q(t)$. Then the local dynamics are governed by the ordinary differential equations
@@
These equations capture the essence of the interaction between the sinusoidal shear profile with amplitude $u$ and turbulent kinetic energy with amplitude $q$. They contain the two most important model closures: $A(q)$ modelling the Reynolds-stress amplitude and $\varepsilon_{0}(q;Re)$ modelling the dominant component of the turbulent pseudo-dissipation rate. (See equations (11), (14), and (15).) The 11 appearing in the expression (1u)(1-u) in momentum equation (24a) comes from the body force that acts against viscous decay and Ekman friction to drive $u$ to toward the laminar value $u=1$. The Reynolds-stress divergence, $\beta A(q)$, is a force acting against the shear whenever $q>0$. Recall that $A(0)=0$. In the TKE equation (24b), TKE is generated by production, $\frac{1}{2}\beta uA(q)$, and dissipated by Ekman friction and the turbulent pseudo-dissipation rate.
The local dynamics described by equations (24) can be understood in terms of nullclines: curves in the $(u,q)$ phase plane on which $\dot{u}=0$ and $\dot{q}=0$. These are shown in figure 6a for two values of ReRe. Steady states of the local dynamics occur where the nullclines intersect and hence $\dot{u}=\dot{q}=0$. We denote steady states by $(u_{ss},q_{ss})$. One steady state is laminar flow $(u_{ss},q_{ss})=(1,0)$ that exists for all ReRe. For ReRe above $Re_{\rm sn}=72.4$, the nullclines intersect at a pair of steady states with $u_{ss}<1$ and $q_{ss}>0$, corresponding to turbulent states. These states are shown in the bifurcation diagram in figure 6b, with turbulent states referred to as upper branch and lower branch depending on the value of $q_{ss}$. These local dynamics are very similar to those in models exhibiting a Turing instability [manneville2012turing, kashyap2025laminar].
@@
Figure 6: (a) Nullclines and (b) bifurcation diagram for local dynamics. In (a), the $u$ nullclines, on which $\dot{u}=0$, are shown in blue and the $q$ nullclines, on which $\dot{q}=0$, are shown in red. Nullclines intersect at steady states of the local dynamics. At $Re=65$, the only steady state is laminar flow $(u_{ss},q_{ss})=(1,0)$, while at $Re=85$, there are two additional steady states. In (b), steady states are plotted as a function of ReRe. $Re_{\rm sn}=72.4$ marks a saddle-node bifurcation in which upper- and lower-branch solutions emerge. Lower-branch solutions are unstable; upper branch solutions are unstable to patterns below $Re_{c}=85.1$ and stable above.
@@ 3.2 Turbulent bands
Going beyond the local dynamics, the model reproduces many features of turbulent bands observed in experiments and DNS of planar shear flows: plane Couette flow, Waleffe flow, and plane Poiseuille flow. We illustrate here some of the most significant dynamics captured by the model. For this we simulate the model equations on a doubly-periodic domain of size $L_{x}\times L_{z}$ using the pseudo-spectral code Dedalus [burns2020dedalus]. Second-order Runge-Kutta time stepping is used with a time step $\Delta t=0.04$, approximately 10 times larger than can be used in a corresponding DNS of the Navier-Stokes equations. The time step in both cases being set based on numerical stability. A Fourier-spectral method with $3/2$ dealiasing is used with a resolution of one grid point per space unit, approximately ten times fewer than for a DNS. Based on the numerical time step and $(x,z)$ spatial discretizations alone, the model is $O(10^{3})$ faster to simulate than the Navier-Stokes equations. Two types of initial conditions (ICs) are used to initialize simulations. A localized IC comprises a finite region of turbulence of length $25$ and width $14$ tilted at $24^{\circ}$ to the $x$ direction. The second IC is uniform turbulence $(u_{0},u_{1},v_{1},w_{0},w_{1},q_{0})(x,z,0)=(0,u_{ss},0,0,0,q_{ss})$. In all cases, the initial large-scale flow fields are additionally seeded with small-scale noise. (To account for turbulent fluctuations on the “microscopic scale”, noise may be further included in simulations. We do not consider this here. The only source of randomness is the ICs.)
@@
Figure 7a shows a simulation at $Re=75$ starting from the localized IC. During the initial evolution, a large-scale, quadrupolar velocity field is established. Such fields are well documented [lundbladh1991direct, schumacher2001largescaleflow, lagha2007largescaleflow, duguet2013oblique, chantry2016turbulent, wang2020quadrupolar, kashyap2020far, couliou2016spreading, klotz2021experimental], and reproducing them is an important validation of the model. The turbulent patch elongates through the growth of its two tips. They eventually join and the system settles into a single steady, straight turbulent band tilted at $\theta=24^{\circ}$ to the streamwise direction [SM, Movies S1-S2]. While the final angle of the band to the streamwise direction is constrained by the domain, during the initial band elongation (figure 7a, $t=750$), the angle is not far from $\theta=24^{\circ}$ (see also figure 9a, $t=2700$, discussed below). Furthermore, doubling the domain size in the spanwise direction still results in $24^{\circ}$ bands [SM, Movie S8]. This is consistent with numerous observations of bands where angles typically line in the range $24^{\circ}\lesssim\theta\lesssim 45^{\circ}$ [tuckerman2020patterns, chantry_universal, prigent2003long, duguet2010formation, duguet2013oblique, ShimizuPRF2019, kashyap2022linear, klotz2022phase, xiong2015turbulent, xiao2020growth]. The large-scale flow directed along the edges of the turbulent bands is typical of that observed in experiments and DNS [coles1966progress, barkley2007mean, duguet2013oblique, gome2023patterns1, marensi2022dynamics]. It is worth emphasizing that the large-scale velocity vectors seen in the midplane are given by $u_{0}(x,z){\bf e}_{x}+w_{0}(x,z){\bf e}_{z}$.
@@
Figure 7b shows a simulation at $Re=78$. At this slightly higher value of ReRe, we observe more growth at the tips and also lateral splitting of the band, eventually resulting in two identical steady bands within the domain [SM, Movies S3-S4]. Figure 7c shows the $Re=78$ case again, but now starting from the uniform IC. Initially, there is competition between symmetrically related orientations (the system is symmetric under spanwise reflections, $z\to-z$). Eventually, one orientation emerges and a single pair of steady bands forms [SM, Movie S5]. Both lateral splitting [barkley2005computational, duguet2010formation, manneville2012growthspot, marensi2022dynamics, lu2022growth] and band competition [prigent2002large, prigent2003long, duguet2010formation, chantry_universal, klotz2022phase, kashyap2020flow, kashyap2022linear] are well-documented phenomena.
@@
Figure 7d shows a vertical slice of the reconstructed turbulence field $q$ (top panel) and large-scale velocities (bottom two panels), taken across the final steady band at $Re=75$ (cyan line in figure 7a, t = 25350). The plots are strikingly similar to vertical slices in DNS of turbulent bands [barkley2007mean, chantry2016turbulent, gome2023patterns1, tuckerman2020patterns]. The top panel highlights the so-called overhang turbulent regions [coles1966progress, barkley2007mean, duguet2013oblique] that are generated through advection of the turbulent field by the mean shear and captured by the $q_{1}$ field in the model. The middle panel visualizes the streamlines of the large-scale flow in the slice plane and the bottom panel shows contours of the well-documented along-band flow (through the slice plane) [barkley2007mean, chantry2016turbulent, tuckerman2014turbulent]. The structure in the wall-normal direction seen in figure 7d highlights that, while model fields are functions of $(x,z)$, they describe non-trivial, three-dimensional flows very much like those seen in DNS.
@@
We now survey the spatiotemporal dynamics of bands in the 1D tilted domain, such that bands are at a fixed angle and independent of the along-band direction. We do so by working in coordinates $({\bar{x}},{\bar{z}})$ rotated by $\theta=24^{\circ}$ with respect to the $(x,z)$ coordinates and setting to zero all derivative in the along-band direction, ${\bar{x}}=x\cos\theta+z\sin\theta$. Restricting to such solutions, the model reduces to a system of 1D PDEs in the across-band coordinate (Appendix A). While restricting the band angle eliminates fully two-dimensional phenomena, such as seen in the early times in figure 7, it provides insights into some of the key spatiotemporal features of turbulent bands.
@@
Figure 8 shows spatiotemporal diagrams at representative values of ReRe, with the band angle fixed at $\theta=24^{\circ}$. Simulations are initiated from a localized IC (a steady band solution at $Re=66$), except in figure 8f where the simulation is initiated with a uniform turbulent IC. At sufficiently low ReRe, turbulent bands are not sustainable and localized initial states rapidly decay, as illustrated at $Re=64$. Note the timescale of panel figure 8a in comparison with the other panels. Starting at $Re\simeq 66$, steady bands are sustained in the form of localized, isolated structures. This is well below the onset of bistability in the homogeneous dynamics in figure 6.
@@
At larger ReRe, figures 8c-8e, an initially localized turbulent patch spreads laterally. At $Re=80$ and $Re=82$, the spreading occurs via what is known as splitting, and the resulting asymptotic states are regular periodic patterns of turbulent bands. (Beyond time shown in figure 8c, the system reaches a steady pattern with eight turbulent bands.) At $Re=90$, turbulence expands rapidly in what is referred to as a slug [wygnanski1973transition, darbyshire1995transition, barkley2015rise]. The asymptotic state is uniform turbulence. As can be seen, the expansion rate is an increasing function of ReRe. The final panel, figure 8e, shows the formation of bands from uniform turbulence at $Re=80$, the same ReRe as in figure 8c. This demonstrates explicitly the instability of the uniform turbulent state at this ReRe.
@@
Figure 8: Spatiotemporal dynamics of 1D bands at a fixed angle $\theta=24^{\circ}$, for various ReRe indicated in the panels. Visualized is the turbulent field $q_{0}(z^{\prime},t)$, where $z^{\prime}$ is the across-band coordinate. (a)-(e) Dynamics starting from a localized turbulent patch illustrating the following behaviour: (a) relaminarization, (b) single, localized turbulent band, (c) and (d) expanding turbulence in the form of band splitting or patterned slug, (e) expanding turbulent slug. (f) Illustration of band formation from uniform turbulence. Panels (c) and (f) correspond to the same conditions, just different initial conditions. Beyond the time shown in (c) the system develops eight turbulent bands.
@@
Finally, returning to the 2D model dynamics, we show in figure 9 some of the model’s behaviour in larger domains—twice the size in each horizontal dimension compared to figure 7—starting from the localized ICs. While the initial evolution (first column of figure 9) resembles that of runs in figure 7, the larger domain better captures the complex process of turbulence proliferation. The single band in figure 9a elongates through the growth of its two tips, as before, but no longer simply reconnects with itself upon reaching the boundaries. Instead, it produces a complex web of chaotic bands [SM, Movie S9]. At slightly higher ReRe, proliferation changes markedly due to band splitting (figure 9b). Turbulence expands in all directions—albeit anisotropically—via the continuous generation of new bands through splitting [SM, Movie S10]. This spreading behaviour is strikingly similar to that observed in plane Couette flow [duguet2010formation, marensi2022dynamics]. At yet higher Reynolds number (figure 9c), the initial spreading occurs more like a front of nearly uniform turbulence, which subsequently develops laminar gaps, resulting in bands [SM, Movie S11]. For all cases, intermediate times reveal long-lived complex dynamics as the system adjusts toward equilibrium. During this phase, bands of different orientations compete (middle column of figure 9). Not all combinations of ICs, Reynolds number, and domain size lead to simple bands. We observe a variety of other steady states, such as criss-crossing bands [SM, Movie S6], non-uniform bands (figure 9b), and instances featuring simple bands of both orientations separated by domain boundaries (figure 9c) [prigent2002large, prigent2003long]. Some runs never reach a steady state: bands nearly form, become unstable, break up, and re-form in a repeating cycle. The dynamics of the model are rich, and there is much more to explore, but we leave this for future work.
@@
Figure 9: Illustrative model dynamics in large domains. Temporal evolution of turbulent bands for three values of ReRe, all initialized with the same localized initial condition. (a) $Re=75$, (b) $Re=78$, and (c) $Re=80$ [SM, Movies S9-S11, respectively]. The colour represents the amplitude of $q_{0}$, the vertically-averaged turbulent kinetic energy.